OpenStudy (anonymous):

medal for whoever helps me!! evaluate the expression; 4^-2 x 2^0

OpenStudy (anonymous):

Do you have 0 as an answer?

OpenStudy (anonymous):

2^0=1

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

its. 8, 16, 1/16, -64

OpenStudy (anonymous):

16.

OpenStudy (anonymous):

4^-2 is same as 1/4^2 which equals 1/16. 1/16 x 1 = 1/16